Jefferson Lab HALL A Logbook for January 10 - All EXCEPT Automatic Entries
Main INDEX
Saturday
309160 01/30/10 15:37 R. Michaels re: helicity in the counting room
Thursday
309159 01/28/10 16:35 A. Camsonne CompSimple run 21914
309152 01/28/10 14:54 mercado Fixing Helicity Connections in Counting House
Wednesday
Tuesday
Monday
Friday
309090 01/22/10 15:40 mercado PREX PMT Setup in R-HRS
Thursday
309074 01/21/10 10:12 ZAFAR AHMED Time differnece. MPS vs. QRT, Helicity and Pair Syn.
Wednesday
309073 01/20/10 15:28 R. Michaels network upgrade completed
Tuesday
309015 01/19/10 10:44 R. Michaels Helicity fibers cleanup (from Riad)
Sunday
309014 01/17/10 12:45 Jack VDC gas system restored
309013 01/17/10 08:09 R. Michaels re: network outage planned for Wed Jan 20
Friday
309011 01/15/10 12:42 ZAFAR AHMED MPS is missing in hall A
309010 01/15/10 07:38 R. Michaels 1-day outage of network on Jan 20
Thursday
309009 01/14/10 16:39 A. Camsonne Octal board added on Serial Controller of Compton IOC
Wednesday
309006 01/13/10 16:35 R. Michaels new CompSimple config for Compton
Friday
309004 01/08/10 15:49 R. Michaels parity daq at 240 Hz
308991 01/08/10 14:29 R. Michaels Helicity flip at 240 Hz now
308990 01/08/10 14:21 brads Counting house is warm and quite humid -- is this 'normal'?
308989 01/08/10 12:52 R. Michaels plan to upgrade network in Hall A (1-day down)
Tuesday
308984 01/05/10 16:48 R. Michaels DAQ mods, PREX
308942 01/05/10 13:01 Kai FADC cables removed
Monday
308941 01/04/10 17:32 A. Camsonne Hit pattern on noise Bob2 configuration run 21817
308939 01/04/10 12:50 Ramesh/Dianc RHRS shower counter TA-51 lemo connector broken (need to fix)
Saturday
308938 01/02/10 15:04 Jack VDC high voltage to remain OFF
Friday
308937 01/01/10 17:36 A. Camsonne Compton ETROC run 21817